OpenStudy (anonymous):

Determine the values of sin 2θ, cos 2θ, and tan 2θ, given tanθ = -7/24, and π/2 ≤ θ ≤ π.

OpenStudy (goformit100):

@apoorvk

OpenStudy (anonymous):

Use them: \[\sin2x = \frac{2tanx}{1 + \tan^2x}\] \[\cos2x = \frac{1 -\tan^2 x}{1 + \tan^2 x}\]

OpenStudy (anonymous):

how did you come up with those ? :S we're supposed to used double angle formulase

OpenStudy (anonymous):

Your sine value and tan value are wrong in denominator I guess @apoorvk

OpenStudy (anonymous):

\[\tan2x = \frac{2tanx}{1 - \tan^2x}\]

OpenStudy (apoorvk):

Sorry for the error!

OpenStudy (anonymous):

It is okay I am just reminding you ..

OpenStudy (anonymous):

are these double angle ?

OpenStudy (anonymous):

Yes they are.. Left hand side you have 2x and right hand side you have x..

OpenStudy (anonymous):

you have given the value of tanx, you have to just plug in that value in place of tanx in the formulas..